Today is July 1, 2000. Youve just borrowed $10,000 to purchase a car. The terms of the loan are 6% APR with monthly compounding Its an exceptionally short term loan- only 6 months total in duration. In addition, you1l only make payments 2, 4, and 6 months from today. The first payment will be $30. The second payment (made 4 months from today) will be $300. Part A) Determine the size of the last payment (made 6 months from today) Parts B/C/D) Complete the amortization schedule given below. Enter the values into PolyLearn which correspond to the three boxes (labeled as Parts B/C/D) below: Principal at End of Month Month # Principal Begin of Month Interest Due Payment Change in Made Principal 0.00 30.00 0.00 300.00 0.00 10,000.00 10,050.00 10,050.00 50.35 -249.40 49.36 Part B 9,871.20 9,920.56 0.00 4 Part C 9,920.56 Part D
Answer #1

Statement showing loan repayment schedule

Month Principal at beginning of month Interest Due @ 0.5% PM Payment made Change in principal Principal at end of month
a b =a*0.5% c d=b-c a+d
1 10000 50 0 50 10050
2 10050 50.25 30 20.25 10070.25
3 10070.25 50.35 0 50.35 10120.60
4 10120.60 50.60 300 -249.40 9871.20
5 9871.20 49.36 0 49.36 9920.56
6 9920.56 49.60 9970.16 -9920.56 0.00